OpenStudy (anonymous):

How do you find the radius of a circle that is inscribed in an isosceles triangle with side lengths 12, 12, and 8?

OpenStudy (anonymous):

SOMEBODY HELP PLEASE!!!

OpenStudy (anonymous):

sorry that was wrong. The formula for radius is r=sqrt[(s-a)(s-b)(s-c)/s] where s=(12+12+8)/2 and a=12 b=12 and c=8. The answer comes to 2.828427